an orifice of area 10sq.mm is made at the bottom of a vessel containing water to a height of 10m above offirice. what is the force on the vessel

Grade:12

1 Answers

Amogh Dasture
7 years ago
Let us assume that mass dm of water flows out in time dt. Let momentum of vessel be P.
 
By Law of conservation of momentum
 
0=P+v(dm)
P= -vdm
 
v is the velocity of efflux which we will assume to be constant
 
F=dP/dt=-v(dm/dt)=-v\rho A(dx/dt)=-\rho Av^{2}
 
Now u can put the value of v2=2gh and get the answer